OpenStudy (anonymous):

Solve this system of equations by using substitution. 3x + y = 1 y = x + 5

OpenStudy (anonymous):

replace \(y\) in the first equation by \(x+5\) and solve \[3x+x+5=1\] for \(x\)

OpenStudy (anonymous):

it has to be two points like this (___,___)

OpenStudy (the_fizicx99):

It will be once you solve it, 3x + x + 5 = 1, solve for x

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (the_fizicx99):

3x + x = ?

OpenStudy (anonymous):

4x

OpenStudy (the_fizicx99):

4x + 5 = 1 Subtract 5 from both sides, whats that?

OpenStudy (anonymous):

4x=-4

OpenStudy (the_fizicx99):

Do you know how to isolate x?

OpenStudy (anonymous):

um like devide both sides

OpenStudy (the_fizicx99):

Yes, 4x = -4 /4 /4 x = ?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

so i know (-1,___)

OpenStudy (the_fizicx99):

Now plug in -1 for x and solve for y, y = -1 + 5

OpenStudy (the_fizicx99):

Solved? I have to leave right now.

OpenStudy (anonymous):

k ty

OpenStudy (the_fizicx99):

yw, you're suppose to get (-1,4) 3(-1) + 4 = 1 -3 + 4 = 1 1 = 1 <-- therefore this works

OpenStudy (anonymous):

yes i got that

OpenStudy (the_fizicx99):

Ok. Bye.